40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  ATF eLog  Not logged in ELOG logo
Entry  Thu Aug 5 09:20:44 2021, Aidan, Laser, 2um Photodiodes, Optical design for 2um PD in new cryo chamber Cryo_Chamber_optical_layout.pdf
    Reply  Thu Aug 5 12:39:56 2021, Aidan, Laser, 2um Photodiodes, Optical design for 2um PD in new cryo chamber Cryo_in-air_setup.pdf
       Reply  Thu Aug 5 14:29:42 2021, Stephen, Laser, 2um Photodiodes, Optical design for 2um PD in new cryo chamber IMG_9685.JPGIMG_9687.JPGrouting_markup_of_IMG_9688.pngIMG_9682.JPG
          Reply  Tue Aug 24 15:01:25 2021, Stephen, Laser, 2um Photodiodes, 2um PD in new cryo chamber 
             Reply  Thu Aug 26 14:31:46 2021, Aidan, Stephen, Laser, 2um Photodiodes, Incident power calibration of 2um PD in new cryo chamber IMG_4420.jpgREF_PD_POW_CAL_cryo.pdf
                Reply  Fri Aug 27 17:57:30 2021, Stephen, DailyProgress, 2um Photodiodes, Start of PD pumpdown  
                   Reply  Sat Aug 28 07:01:21 2021, Aidan, DailyProgress, 2um Photodiodes, Start of PD pumpdown  20210828_-_IMG_4445.jpg20210828_-_IMG_4443.jpg
                      Reply  Sun Aug 29 11:17:43 2021, Aidan, DailyProgress, 2um Photodiodes, Start of PD pumpdown  Screen_Shot_2021-08-29_at_11.16.55_AM.pngScreen_Shot_2021-08-29_at_11.22.03_AM.png
                         Reply  Mon Aug 30 08:04:52 2021, Aidan, DailyProgress, 2um Photodiodes, Monday morning observation of JPL A1 PD cool down  Screen_Shot_2021-08-30_at_8.04.15_AM.png
Message ID: 2650     Entry time: Tue Aug 24 15:01:25 2021     In reply to: 2639     Reply to this: 2651
Author: Stephen 
Type: Laser 
Category: 2um Photodiodes 
Subject: 2um PD in new cryo chamber 

[Aidan, Stephen]

Worked toward aligning and characterizing beam on PD. Will complete next session.

Log:

Some difficulty aligning to the 2um beam, which is sensed by a thermal card. Aidan intends to upgrade with a fiber coupled visible laser, which could then be swapped interchangably for alignment.

The 1" mirror at the top of the periscope doesn't make sense, given larger apertures in shields and viewport. We looked for a nearby 2" replacement but did not have luck. Ended up swapping back in the gold-coated 2" mirror, even though it is thin enough to be a pain to mount.

Instrumented connector pins to DB9 pins using the following translation (ref Aidan's drawing for connector / PD pinout, ref drawing from QIL/2639)

            DB9   - 1 6 2 7 3 8 

    Connector - 2 3 4 5 6 7

 

ELOG V3.1.3-